# Experimental Demonstration of a Decentralized Electromagnetic Formation Flying Control Using Alternating Magnetic Field Forces

### v1

@@ 5.1 Desired Intersatellite Forces
If $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$, then it follows from 7 the approximate average intersatellite force is
@@
In this case, [1, Theorem 5] shows how to select the gains $\alpha_{ij}$ and $\beta$ such that the formation error $\mathbf{r}_{ij,k}-\mathbf{d}_{ij}$ tends to zero for $\hat{\mathbf{F}}_{ij,k}=\hat{\mathbf{F}}_{ij,k}$. The approximate average intersatellite force 9 is based on linear consensus for sampled-data double integrators [5]. This method uses feedback to determine approximate average intersatellite forces $\hat{\mathbf{F}}_{ij,k}$ that create virtual linear springs and dashpots between satellite $i$ and $j$.
@@ 5.2 Control Amplitude Pair To Achieve Prescribed Intersatellite Force
This section provides a construction for the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ such that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$. This construction is presented in [9, 8] for the system resolved in an inertial frame. Here, we present the construction coordinate free.
Define
@@
Next, define
$\mathbf{h}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq\begin{cases}\frac{h_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}\mathbf{r}+\frac{h_{\mathrm{rf}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}||\mathbf{r}\times\mathbf{f}^{*}|}((\mathbf{r}\times\mathbf{f}^{*})\times\mathbf{r}),&\mathbf{r}\times\mathbf{f}^{*}\neq\mathbf{0},\\ \frac{h_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}\mathbf{r},&\mathbf{r}\times\mathbf{f}^{*}=\mathbf{0}.\end{cases}$ (17)
The following result shows that 16 and 17 is a construction for the amplitude pair such that $\mathbf{f}$ takes on a prescribed value $\mathbf{f}^{*}$. The proof is in A.
@@
Proposition 2 implies that we can use 16 and 17 to select the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ such that the average intersatellite force equals the desired value 8. Specifically, we let the amplitudes be
@@
and it follows from Proposition 2 that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$. In other words, the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ given by (18) achieves the desired intersatellite force $\mathbf{f}^{*}_{ij,k}$.
@@ 6 Experiment Setup
This section describes a one-dimensional (1D) experimental platform for testing the EMFF algorithm from Sections 4 and 5, specializes the algorithm to 1D, and presents modifications for implementation.
@@ 6.1 Experimental Platform
Each EAS unit has an electromagnetic coil and an electrical and electronics platform. Each electromagnetic coil has an area of $A=0.1\pi^{2}$ m2 and consists of $N=500$ turns of 22 gauge copper wire. The electrical and electronics platforms include all components needed to supply power to the electromagnetic coils. The power electronics consist of two lithium-ion polymer batteries, two Vicor DC-DC converters, and a Copley Controls Junus JSP-090-10 amplifier, as shown in Figure 3.
Each EAS unit has an Arduino Due microcontroller which is responsible for data acquisition and feedback control. An XBee radio frequency (RF) module is connected to an Arduino wireless secure digital (SD) shield, which is attached to the microcontroller. A coordinator Xbee RF module is also connected to a computer, enabling communication between the computer and the EAS units through a Digi XCTU application.
@@ 6.2 Control in 1D with Current Implementation
where
@@
Substituting 23, 24, and 25 into 27 yields $f(r_{ij,k},p_{ij,k},p_{ji,k})=f^{*}_{ij,k}$, which combined with 20 and 26 demonstrates that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$.
@@ 6.3 Control Modification to Address Hardware Limitations
We introduce 2 modifications to the current control 21, 22, 23, and 24. First, we introduce integral action to the desired force 21 to address steady-state error. Second, we introduce modifications to the current allocation 23 to address current amplitude saturation. This section redefines $f_{ij,k}^{*}$ and $I_{ij,k}$ given by 21 and 23 based on these modifications.
@@
where $\xi_{ij,0}\in\mathbb{R}$ is the initial condition. Note that 28 is an integrator if the formation error $r_{ij,k}-d_{ij}$ has small magnitude (i.e., in the interval $(\epsilon_{0},\epsilon_{1})$). Otherwise, the integrator is disabled. Then, define the modified desired formation force
@@
which is the maximum amplitude of the sum of sinusoidal currents over $t\in[kT,kT+T)$. Since the current is limited by the power capability of the EAS units, we modify $I_{ij,k}$ such that $\bar{I}_{i,k}$ does not exceed the maximum allowable current $\bar{I}>0$. The current implemented on the $i$th coil is 22, where the current amplitude is
@@
Then, it follows from 22 and 30 that
@@
Next, we show that the resultant prescribed force is a scaling of the unconstrained prescribed force $f^{*}_{ij,k}$. Specifically, substituting 24, 25, and 30 into 27 yields
@@
Furthermore, 6 and 7 imply that the approximate average intersatellite force generated is equal to $\hat{\mathbf{F}}_{ij,k}=\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})\mathbf{i}$, where
@@
6.4 Estimation of Relative Position and Relative Velocity
@@
$-\mathopen{}\mathclose{{\left.\sum_{h\in\mathcal{N}_{j}}\hat{F}(\hat{r}_{jh,k},p_{jh,k},p_{hj,k})}}\right).$ (33)
@@
$-\mathopen{}\mathclose{{\left.\sum_{h\in\{i\}\cup(\mathcal{N}_{i}\cap\mathcal{N}_{j})}\hat{F}(\hat{r}_{jh,k},p_{jh,k},p_{hj,k})}}\right).$ (34)
@@
$\begin{bmatrix}\hat{r}_{ij,k}\\ \hat{v}_{ij,k}\end{bmatrix}$= $(A-LCA)\begin{bmatrix}\hat{r}_{ij,k-1}\\ \hat{v}_{ij,k-1}\end{bmatrix}+(B-LCB)\hat{\nu}_{ij,k-1}$
$+Lq_{ij,k},$ (35)
@@
where $P$ is the solution to the discrete algeabric Ricatti equation
@@
where $W\triangleq wBB^{\mathrm{T}}$ and $w>0$ is the variance of the disturbance force.
@@ 7 Experimental Results and Discussion
We present results from open-loop and closed-loop experiments using the EAS units described in Section 6.1 with the control algorithm presented in Section 6.2 with the modifications presented in Section 6.3. The plots of the experiments include $|q_{ij,k}|$, $|\hat{r}_{ij,k}|$, $\hat{v}_{ij,k}$, $I_{ij,k}$, and $\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})$. The Kalman filter is initialized with $\hat{r}_{ij,0}=q_{ij,0}$ and $\hat{v}_{ij,0}=0$ m/s.
7.1 2 Satellite Experiments
We present 2 open-loop experiments and 3 closed-loop experiments with 2 satellites. The left satellite is numbered 1 and the right satellite is numbered 2. The actuation frequency are $\omega_{12}=\omega_{21}=40\pi$ rad/s, and the control is implemented at $T=0.1$ s. Using experimental data, we estimate the variances $V=1.2\times 10^{-6}$ $\mathrm{m}^{2}$ and $w=5\times 10^{-6}$ $\mathrm{m}^{2}/\mathrm{s}^{4}$. It follows from 36, 37, and 38 that $L=\begin{bmatrix}0.0942&0.0466\end{bmatrix}^{\mathrm{T}}.$
@@
This experiment demonstrates open-loop attraction. Figure 4 shows the experimental results, where the first row shows the raw relative-position measurements $|q_{ij}|$, the second and third rows show the filtered measurements $|\hat{r}_{ij}|$ and $\hat{v}_{ij}$, the fourth row shows the current $I_{ij}$, and the fifth row shows the average intersatellite force $\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})$. The first column shows the data measured and computed onboard satellite 1, and the second column shows the data measured and computed onboard satellite 2.
@@
Next, we present 3 experiments demonstrating closed-loop control. The control $(I_{12,k},I_{21,k})$ is implemented with $\alpha_{ij}=0.0158$, $\beta=6.89$, $\rho_{ij}=0$, and $\bar{I}=2.35$ A. Note that $\alpha_{ij}$ and $\beta$ are selected such that $\alpha_{ij}m/(c_{0}N^{2}A^{2})=812$ and $\alpha_{ij}\beta m/(c_{0}N^{2}A^{2})=5600$.
@@
This experiment demonstrates closed-loop repulsion. The satellites start at a relative distance of approximately 0.4 m. The objective is to reach a desired relative distance $|d_{ij}|=0.45$ m. Figure 6 shows that $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.45$ m after approximately 20 s, which demonstrates that the satellites achieve formation. There is a sharp decrease in the measurement $|q_{12}|$ at $t=2.5$ s, which is due to a sensor dropout.
@@
$P\triangleq\sqrt{\frac{1}{N_{\mathrm{t}}}\sum_{k=0}^{N_{\mathrm{t}}}\mathopen{}\mathclose{{\left(\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})}}\right)^{2}},$ (39)

### v2

@@ 7.1 Desired Intersatellite Forces
If $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$, then it follows from 7 that
@@
In this case, [1, Theorem 5] provides conditions on selecting the gains $\alpha_{ij}$ and $\beta$ such that the formation error $\mathbf{r}_{ij,k}-\mathbf{d}_{ij}$ tends to zero for $\bar{\mathbf{F}}_{ij}(k)=\hat{\mathbf{F}}_{ij,k}$. The approximate time-averaged intersatellite force 9 is based on linear consensus for sampled-data double integrators [6]. This method uses feedback to determine approximate time-averaged intersatellite forces $\hat{\mathbf{F}}_{ij,k}$ that create virtual linear springs and dashpots between satellite $i$ and $j$. We note that larger $\alpha_{ij}$ tends to result in faster formation convergence but can cause oscillations (e.g., overshoot) in the response. Larger $\beta$ increases the virtual intersatellite damping and tends to reduce oscillations. We also note that the analysis in [1, Theorem 5] relies on the assumption that $\bar{\mathbf{F}}_{ij}(k)=\hat{\mathbf{F}}_{ij,k}$, which is a reasonable approximation if $\mathbf{r}_{ij}$ does not change significantly over $[kT,kT+T)$.
@@ 7.2 Control Amplitude Pair To Achieve Prescribed Intersatellite Force
This section provides a construction for the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ such that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$. This amplitude pair is constructed from $\mathbf{r}_{ij,k}$ and $\mathbf{f}^{*}_{ij,k}$. The construction is presented in [9, 10] for physical vectors resolved in an inertial frame. Here, we present the construction coordinate free. The following functions of a relative position $\mathbf{r}$ and desired force function value $\mathbf{f}^{*}$ are needed to construct the control amplitudes. Specifically, we define
$\mathbf{h}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq\begin{cases}\frac{h_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}\mathbf{r}+\frac{h_{\mathrm{rf}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}||\mathbf{r}\times\mathbf{f}^{*}|}((\mathbf{r}\times\mathbf{f}^{*})\times\mathbf{r}),&\mathbf{r}\times\mathbf{f}^{*}\neq\mathbf{0},\\ \frac{h_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}\mathbf{r},&\mathbf{r}\times\mathbf{f}^{*}=\mathbf{0},\end{cases}$ (11)
where
@@
The following result shows that 10 and 11 are an amplitude pair such that $\mathbf{f}$ takes on the prescribed value $\mathbf{f}^{*}$.
@@
Proposition 2 shows how to select the amplitudes $\mathbf{p}_{ij,k}$ and $\mathbf{p}_{ji,k}$ of the common frequency sinusoids in 4. In other words, Proposition 2 implies that we can use 10 and 11 to select $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ such that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$, which implies that the approximate time-averaged intersatellite force equals the desired value given by 9. Specifically, we let the amplitudes be
@@
It follows from Proposition 2 that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$. In other words, the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ given by (18) achieves the desired value of the intersatellite force function 8.
First, consider (i) $\mathbf{r}\times\mathbf{f}_{*}\neq\mathbf{0}$. Substituting $|\mathbf{r}\times\mathbf{f}_{*}|=|\mathbf{r}|^{2}\mathbf{f}_{*}-(\mathbf{r}\cdot\mathbf{f}_{*})\mathbf{r}$, (10), and (11) into (2) yields
$g_{\mathrm{r}}h_{\mathrm{rf}}$ $=g_{\mathrm{rf}}h_{\mathrm{r}}=\frac{|\mathbf{r}\times\mathbf{f}_{*}|}{2|\mathbf{r}|},$
$a_{y}b_{y}$ $=-\frac{\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}_{*})}{2|\mathbf{r}|}\mathopen{}\mathclose{{\left(-|\mathbf{r}\cdot\mathbf{f}_{*}|+\Phi_{1}}}\right).$
Next, consider (ii) $\mathbf{r}\times\mathbf{f}_{*}=\mathbf{0}$, and substituting (10)–(15) into (2) yields
@@ 8 Experiment Setup
This section describes a one-dimensional (1D) experimental platform for testing the EMFF algorithm from Sections 6 and 7, specializes the algorithm to 1D, and presents modifications for implementation.
@@ 8.1 Testbed Hardware
Each EAS unit has an electromagnetic coil and an electrical and electronics platform. Each electromagnetic coil has an area of $A=0.1\pi^{2}$ m2 and consists of $N=500$ turns of 22 gauge copper wire. The impedance of each coil is 71 mH and the resistance is 16 Ohms. The electrical and electronics platforms include all components needed to supply power to the electromagnetic coils. The power electronics consist of two lithium-ion polymer batteries, two Vicor DC-DC converters, and a Copley Controls Junus JSP-090-10 amplifier, as shown in Figure 4. Bench tests were performed to confirm that the amplifiers can produce sinusoidal current commands with the frequency content used in the experiments (i.e., 10 Hz and 20 Hz).
Each EAS unit has an Arduino Due microcontroller which is responsible for data acquisition and feedback control. An XBee radio frequency (RF) module is connected to an Arduino wireless secure digital (SD) shield, which is attached to the microcontroller. A coordinator XBee RF module is also connected to a computer, enabling communication between the computer and the EAS units through a Digi XCTU application. The baud rate of the XBees are set at 115,200, and the data transmission frequency is 2.4 GHz. The satellites’ Arduinos start the experiment simultaneously using a transmission from the XBee coordinator. The control is updated with period $T$ to limit frequency drift between the satellites. Although there was no significant time or frequency drift in the experiments, this could be a concern for long-duration maneuvers, necessitating periodic resynchronization of the satellites or adaptive frequency tuning.
@@ 8.2 Control in 1D with Current Implementation
where the desired value for the intersatellite force function in the $\mathbf{i}$ direction is
@@
Substituting 25, 26, and 27 into 29 yields $f(r_{ij,k},p_{ij,k},p_{ji,k})=f^{*}_{ij,k}$, which combined with $2$ and 28 demonstrates that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$.
@@ 8.3 Control Modification to Address Hardware Limitations
We introduce 2 modifications to the current control 23, 24, 25, and 26. First, we introduce integral action to the desired force 23 to address steady-state error. Second, we introduce modifications to the current allocation 25 to address current amplitude saturation. This section redefines $f_{ij,k}^{*}$ and $I_{ij,k}$ given by 23 and 25 based on these modifications.
@@
where $\xi_{ij,0}\in\mathbb{R}$ is the initial condition. Note that 30 is an integrator if the formation error $r_{ij,k}-d_{ij}$ has small magnitude (i.e., in the interval $(\epsilon_{0},\epsilon_{1})$). Otherwise, the integrator is disabled to prevent integrator windup during the transient phase when the position error can be relatively large. Then, define the modified desired formation force
@@
which is the maximum amplitude of the sum of sinusoidal currents over $t\in[kT,kT+T)$. Since the current is limited by the power capability of the EAS units, we modify $I_{ij,k}$ such that $\bar{I}_{i,k}$ does not exceed the maximum allowable current $\bar{I}>0$. The current implemented on the $i$th coil is 24, where the current amplitude is
@@
Then, it follows from 24 and 32 that
@@
Next, we show that the resultant prescribed force is a scaling of the unconstrained prescribed force $f^{*}_{ij,k}$. Specifically, substituting 26, 27, and 32 into 29 yields
@@
Furthermore, 6 and 7 imply that the approximate time-averaged intersatellite force is $\hat{\mathbf{F}}_{ij,k}=\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})\mathbf{i}$, where
@@
8.4 Relative Position and Velocity Estimation Using a Kalman Filter
@@
$\qquad-\mathopen{}\mathclose{{\left.\sum_{h\in\mathcal{N}_{j}}\hat{F}(\hat{r}_{jh,k},p_{jh,k},p_{hj,k})}}\right).$ (35)
@@
$\qquad-\mathopen{}\mathclose{{\left.\sum_{h\in\{i\}\cup(\mathcal{N}_{i}\cap\mathcal{N}_{j})}\hat{F}(\hat{r}_{jh,k},p_{jh,k},p_{hj,k})}}\right).$ (36)
@@
$\begin{bmatrix}\hat{r}_{ij,k}\\ \hat{v}_{ij,k}\end{bmatrix}=(A-LCA)\begin{bmatrix}\hat{r}_{ij,k-1}\\ \hat{v}_{ij,k-1}\end{bmatrix}+(B-LCB)\hat{\nu}_{ij,k-1}+Lq_{ij,k},$ (37)
@@
where $P$ is the solution to the discrete algebraic Riccati equation
@@
where $W\triangleq wBB^{\mathrm{T}}$ and $w>0$ is the variance of the disturbance force. The performance of this filter may degrade in the presence of non-Gaussian disturbances (e.g., disturbances from the airflow leaving the airtracks in the experiments, atmospheric drag in low Earth orbit for on-orbit deployment). In this case, other state estimation methods (e.g., particle filters) could be considered.
@@ 9 Experimental Results and Discussion
We present results from open-loop and closed-loop experiments using the EAS units described in Section 8.1 with the control algorithm presented in Section 8.2 with the modifications presented in Section 8.3. The control update period $T=0.1$ s is selected based on a tradeoff between the computational capability of the microprocessor and the speed of the closed-loop response. Recall that the approximate time-averaged formation dynamics (i.e., (5) with $\bar{\mathbf{F}}_{ij}(k)=\hat{\mathbf{F}}_{ij,k}$ is developed under the assumption that $\mathbf{r}_{ij}$ does not change significantly over period $T$. Thus, $T$ is selected to be small enough that this assumption is reasonable and large enough such that all computations can be performed in real time onboard the microprocessor. The plots of the experiments include $|q_{ij,k}|$, $|\hat{r}_{ij,k}|$, $\hat{v}_{ij,k}$, $I_{ij,k}$, and $\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})$. The Kalman filter is initialized with $\hat{r}_{ij,0}=q_{ij,0}$ and $\hat{v}_{ij,0}=0$ m/s.
9.1 Two-Satellite Validation in Open and Closed Loop
We present 2 open-loop experiments and 3 closed-loop experiments with 2 satellites. The left satellite is numbered 1 and the right satellite is numbered 2. The actuation frequency are $\omega_{12}=\omega_{21}=40\pi$ rad/s. Using experimental data, we estimate the variances $V=1.2\times 10^{-6}$ $\mathrm{m}^{2}$ and $w=5\times 10^{-6}$ $\mathrm{m}^{2}/\mathrm{s}^{4}$. It follows from 38, 39, and 40 that
${P=\begin{bmatrix}0.2686&0.2710\\ 0.2710&0.5206\\ \end{bmatrix}\times 10^{-6}},\qquad L=\begin{bmatrix}0.2013\\ 0.1845\end{bmatrix}$.
@@ 9.1 Two-Satellite Validation in Open and Closed Loop
This experiment demonstrates open-loop attraction. Figure 5 shows the experimental results, where the first row shows the raw relative-position measurements $|q_{ij}|$, the second and third rows show the filtered measurements $|\hat{r}_{ij}|$ and $\hat{v}_{ij}$, the fourth row shows the current $I_{ij}$, and the fifth row shows the time-averaged intersatellite force $\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})$. The first column shows the data measured and computed onboard satellite 1, and the second column shows the data measured and computed onboard satellite 2.
@@
Next, we present 3 experiments demonstrating closed-loop control. The control $(I_{12,k},I_{21,k})$ is implemented with $\alpha_{ij}=0.0158$, $\beta=6.89$, $\rho_{ij}=0$, and $\bar{I}=2.35$ A. Note that $\alpha_{ij}$ and $\beta$ are selected such that $\alpha_{ij}m/(c_{0}N^{2}A^{2})=812$ and $\alpha_{ij}\beta m/(c_{0}N^{2}A^{2})=5600$. These gains are selected based on a tradeoff between response speed and hardware capability (e.g., current limits). The selected values are such that the settling time and overshoot for the idealized closed-loop double-integrator dynamics are approximately 20 s and less than 0.01 m, respectively.
@@
This experiment demonstrates closed-loop repulsion. The satellites start at a relative distance of approximately 0.4 m. The objective is to reach a desired relative distance $|d_{ij}|=0.45$ m. Figure 7 shows that $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.45$ m after approximately 20 s. Specifically, the settling time is $21.1$ s, and the mean steady-state error $\hat{r}_{ij}-d_{ij}$ over the last 60 s is $2.7\times 10^{-3}$ m. Note that there is a sharp decrease in the measurement $|q_{12}|$ at $t=2.5$ s, which is due to a sensor dropout. The variance of the unfiltered steady-state error $q_{ij}-d_{ij}$ is $3.1\times 10^{-6}$ m2, where $1.2\times 10^{-6}$ m2 is attributed to sensor noise and the remaining is potentially due in part to unmodeled external disturbances.
@@
$P_{\hat{F}}\triangleq\sqrt{\frac{1}{N_{\mathrm{t}}}\sum_{k=1}^{N_{\mathrm{t}}}\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})^{2}},$ (41)